# AraFinNews: Arabic Financial Summarisation with Domain-Adapted LLMs

### v2

@@ V-A Arabic Financial Headline Generation
All models were fine-tuned on AraFinNews using identical training settings for fair comparison. Specifically, we fine-tuned:
  • AraT5 (small, base) [nagoudi_arat5_2022]: a monolingual Arabic text-to-text transformer trained on diverse general-domain corpora.
@@
  • FinAraT5 (base): a domain-adapted Arabic T5 variant further pretrained on Modern Standard Arabic financial corpora distinct from AraFinNews.
  • BERT2BERT [rothe-etal-2020-leveraging]: a multilingual encoder–decoder model using mBERT for both encoding and decoding, included as a generic non-T5 generative baseline.
  • Lead-1: an extractive lower-bound baseline that selects the first sentence of each article as its summary.
@@ V-B Evaluation
Model performance was evaluated using ROUGE [lin_rouge_2004], METEOR [banerjee-lavie-2005-meteor], and a suite of embedding-based metrics including BERTScore (BertS), BARTScore (BartS), FrugalScore [kamal-eddine-etal-2022-frugalscore], and BLEURT [sellam2020bleurt]. ROUGE and METEOR capture lexical and shallow semantic overlap, whereas BERTScore, BARTScore, and FrugalScore quantify embedding-based similarity. BLEURT provides a learned measure of coherence.
@@
Overall, AraFinNews offers a valuable benchmark for assessing and improving Arabic financial text generation, promoting more accurate, concise, and domain-aware summarisation systems.
@@ VI Conclusion
This paper introduced AraFinNews, a large-scale Arabic financial news dataset, and demonstrated its use in domain-specific abstractive summarisation through FinAraT5—a Financial Arabic Large Language Model derived from AraT5 via continued pretraining on financial corpora. With over 200,000 article–headline pairs, AraFinNews constitutes the largest publicly available dataset of its kind, providing a valuable benchmark for downstream financial NLP applications such as sentiment analysis, event detection, and question answering.
Experimental results showed that FinAraT5 consistently outperforms both monolingual and multilingual baselines across all automatic metrics, achieving improvements of up to +3.1 in ROUGE-L, +2.4 in METEOR, and +2.8% in FrugalScore over AraT5 and mT5. These gains highlight the effectiveness of domain-adaptive pretraining in producing summaries that are stylistically aligned with Arabic financial journalism. The analysis of attention distributions further demonstrated FinAraT5’s capacity to model fine-grained semantic and numerical relationships characteristic of financial language.

### v3

@@ V-A Arabic Financial Headline Generation
All models were fine-tuned on the AraFinNews training split using identical training settings for fair comparison. Specifically, we fine-tuned:
  • AraT5 (small, base) [nagoudi_arat5_2022]: a monolingual Arabic text-to-text transformer trained on general-domain corpora.
@@
  • FinAraT5 (base): a domain-adapted Arabic T5 variant pretrained on Modern Standard Arabic financial text.
  • BERT2BERT [rothe-etal-2020-leveraging]: a multilingual encoder–decoder model using mBERT for both encoding and decoding.
  • Lead-1: an extractive lower-bound baseline that selects the first sentence of each article as the summary.
@@ V-B Evaluation
Model performance was evaluated using ROUGE [lin_rouge_2004], METEOR [banerjee-lavie-2005-meteor], and a suite of embedding-based metrics including BERTScore (BertS), BARTScore (BartS), FrugalScore [kamal-eddine-etal-2022-frugalscore], and BLEURT [sellam2020bleurt]. ROUGE and METEOR capture lexical and shallow semantic overlap, whereas BERTScore, BARTScore, and FrugalScore (Frugal-1 and Frugal-2) quantify embedding-based similarity. BLEURT provides a learned measure of coherence.
@@
Taken together, these results highlight the clear benefits of domain-adapted modelling for Arabic financial text. AraFinNews provides a robust benchmark for advancing this line of work, supporting the development of more accurate, concise, and context-aware summarisation systems for professional Arabic financial communication.
@@ VI Conclusion
This paper introduced AraFinNews, the largest publicly available Arabic financial news dataset, comprising 212,500 article–headline pairs collected across a decade of reporting. AraFinNews provides a robust benchmark for evaluating financial-domain text generation and supports a range of downstream applications, including sentiment analysis, event detection, and question answering.
We evaluated several monolingual, multilingual, and domain-adapted transformer models on abstractive financial headline generation. Across all automatic metrics, FinAraT5 delivered the strongest performance, achieving consistent gains over general-purpose baselines. These results underline the value of domain-adaptive pretraining for modelling the stylistic and numerical characteristics of Arabic financial journalism. AraFinNews offers both a challenging benchmark and a foundation for advancing Arabic financial NLP.